A merchant purchased a line of handbags for R172. 00 each. She wants to offer a 14% discount off the listed selling price and still make a 24% profit. The price that she should mark on the label is


1
Expert's answer
2021-04-26T15:02:13-0400

The merchant spent R172.00 handbags.

24% of this price is

"\frac{24}{100}\cdot 17200=4128"

So, he has to sell the bag at 17200+4128=21328

14% - discount: 100%-14%=86%

 The label price is such that

"\frac{86}{100}\cdot x=21328\\
x=21328: 0.86\\
x=24800"

The price that she should mark on the label is 24800